# A Perspective on Quantum Computing Applications in Quantum Chemistry using 25–100 Logical Qubits

### v1

@@ III.2 Quantum Dynamics and Noise-Informed Simulation
Quantum devices may be well-suited to this regime due to their natural ability to implement unitary time evolution and sample from high-dimensional entangled states. Methods such as Trotterized real-time evolution, variational dynamics [113], and Krylov-subspace propagation (related to quantum signal processing, or Qubitization) have been proposed as viable quantum algorithms (see e.g. Ref. 107 for a recent overview). However, challenges remain, including the accurate preparation of initial states (which can be non-trivial and propagate errors), significant measurement overhead due to wavefunction collapse, and the difficulty of accurate Trotterization, especially for coupled, time-dependent systems.
@@
Simulating OQS presents the additional challenge that quantum computers naturally perform unitary evolution, while open systems exhibit non-unitary dynamics due to environmental interaction. Standard quantum approaches to OQS include embedding the system within a larger environment simulated unitarily (ancilla-based methods), using stochastic quantum trajectories, or implementing Kraus operators [114, 115, 116, 117, 118, 119, 120]. Intriguingly, noise, traditionally viewed as an impediment, has been proposed as a resource [121]. Hardware-induced decoherence, if properly characterized or engineered, may serve as a proxy for environmental interactions, thereby facilitating OQS simulation. This “noise-assisted” approach has been explored in analog and digital quantum computing [122, 123, 124, 125], potentially reducing overhead compared to full error mitigation by using techniques like partial error correction [126, 127] or pulse control.
@@
The encoding of bosonic modes—representing vibrational, solvent, or bath degrees of freedom—also requires careful consideration. Strategies include truncated Fock spaces, coherent-state encodings, and squeezed-state representations [128]. Opportunities arise from the potentially simpler Hamiltonian structures in exciton-boson systems ( or terms) compared to standard electronic structure ( terms). Furthermore, embedding approaches and the development of advanced Gaussian ansatzes are being explored to reduce the resource requirements for representing these bosonic degrees of freedom [129]. Non-Hermitian Hamiltonians are also relevant for modeling dissipation, although simulating their dynamics on quantum hardware often involves mapping back to larger unitary systems or specific simulation techniques [119].
@@ III.3 Hybrid Pipelines and AI Integration
The integration of quantum computing (QC) with artificial intelligence (AI) and high-performance computing (HPC) is rapidly advancing, with major initiatives developing platform-level solutions for accelerating chemistry and materials discovery [130, 131, 132, 47, 133, 134]. At the forefront of innovation is NVIDIA’s DGX Quantum system, which enables low-latency, tightly coupled execution between quantum processing units (QPUs) and GPUs to support real-time AI-assisted quantum error correction (QEC), calibration, control, and readout [130]. To program such heterogeneous quantum–classical systems, NVIDIA introduced the CUDA-Q platform [135], a single-source, hardware-agnostic programming framework that unifies quantum and classical workflows, leveraging NVIDIA’s existing CUDA and AI ecosystems. Another approach is provided by Microsoft’s Discovery platform (formerly, Azure Quantum Elements), which offers a cloud-based solution that integrates access to multiple quantum hardware backends with Azure’s HPC infrastructure and specialized AI tools, enabling users to build scalable hybrid quantum applications within a managed cloud environment [131].
In such hybrid ecosystems, AI models—often trained on large synthetic datasets generated via HPC simulations or augmented with quantum data—are used to accelerate discovery by enabling rapid screening, property prediction, and system-level optimization. For example, AI models deployed within Azure Quantum Elements have been used to evaluate millions of potential battery materials [136], while NVIDIA has developed generative AI approaches such as GQE [137] and QAOA-GPT [138] to automate the synthesis of optimal quantum circuits with desirable characteristics, such as low depth and high expressivity. In addition to algorithm development, NVIDIA has also demonstrated the use of AI for QEC in collaboration with QuEra [139].
This tiered, adaptive computational approach allocates quantum resources to the most challenging subproblems where classical methods struggle—such as transition states poorly captured by DFT or systems with strong multireference character [140, 141]. AI is integral across this hybrid workflow: from hardware design and calibration to device control, algorithm optimization, QEC decoding, and post-processing [142, 143]. Diagnostic tools—including uncertainty quantification in AI models, convergence analysis of classical solvers (e.g., DMRG), and AI-driven screening and circuit optimization—support dynamic resource allocation across the classical-quantum-AI stack [144, 145, 146].
@@
AI’s role spans the entire quantum computing stack [130]. In hardware development, it accelerates qubit characterization, architecture exploration, and control pulse optimization. During operation, AI automates calibration and tuning, enabling closed-loop control strategies adaptive to evolving noise environments. In software, AI aids circuit synthesis and compression, variational optimization, and hybrid workload scheduling. It enhances QEC by improving decoder performance [29] and enabling scalable, low-latency strategies. In post-processing, AI reduces measurement overhead and mitigates errors in tasks like tomography [147], observable estimation, and readout classification. This end-to-end integration of AI is essential for scalable, fault-tolerant quantum computing and efficient use of hybrid resources.
@@ III.4 Benchmarking and Experimental Validation
Robust benchmarking remains essential for validating quantum methods and ensuring reproducibility. A three-way validation framework—incorporating quantum simulations, high-level classical reference data, and experimental observations—has been endorsed [43, 44, 21]. This framework allows for iterative refinement of algorithms and facilitates the identification of performance gaps.
Benchmark systems that are well-defined chemically relevant, experimentally tractable, and computationally challenging have been proposed as community-wide testbeds. These systems allow simulations at different levels and experiments to interact closely, enabling to guide method development, hardware requirements, and software stacks across multiple disciplines. Table 1 summarizes several classes of model systems identified as focal points for algorithm validation and workflow integration. These span catalysis, photochemistry, and energy materials.
@@ IV Algorithmic Innovations
The development of quantum algorithms capable of addressing chemically relevant problems using 25–100 logical qubits must be guided by co-design principles that account for fault-tolerant constraints. These include limited circuit depth, restricted qubit connectivity, noise resilience, and modular architecture [167, 168]. The goal is to create a unified framework for evaluating algorithms across diverse problems, establishing robust metrics and methodologies to uniformly measure performance [169]. Rather than converging on a single dominant paradigm, the current landscape favors the parallel exploration of diverse algorithmic strategies tailored to specific problem structures and hardware capabilities [142, 20, 21]. Many algorithms exist, balancing accuracy and efficiency, as exact solutions scale exponentially and are often impractical for large systems. Below, we highlight several promising algorithmic directions in this context.
IV.1 Structured Ansatzes and Adaptable Methods
Traditional variational quantum eigensolver (VQE) approaches [61] employing unitary coupled-cluster (UCC) ansatzes face challenges such as barren plateaus, spin contamination, and instability across geometries [26]. Developments in areas like neural network ansätze for solids [170], most innovation has been focused on more chemically inspired and hardware-efficient ansatzes [32, 171]:
@@ IV.1 Structured Ansatzes and Adaptable Methods
Real, imaginary, and complex parameterizations of the unitary Cluster Jastrow (uCJ) ansatz [172, 173]. For instance, the real antisymmetric uCJ form is widely used and offers a more physically meaningful treatment of strong correlation than the generalized unitary coupled-cluster with singles and doubles (GUCCSD) [174], while significantly reducing quantum resource requirements [173]. Alternative parameterizations—such as the imaginary symmetric (Im-uCJ) and general complex (g-uCJ) forms—are also promising, particularly in challenging scenarios like bond dissociation. These findings suggest that tuning the parameterization strategy can improve accuracy and robustness, though care must be taken to evaluate potential drawbacks, such as spin contamination.
@@
Optimization-free and adjustable subspace construction methods. Representative examples include quantum subspace expansion [175], non-orthogonal configuration interaction approaches [167], hybrid and quantum Krylov/Lanczos methods [176, 177, 178, 179], quantum computed moment techniques [180], the quantum equation-of-motion approach [181, 182], and generator coordinate-inspired strategies [183]. The subspace basis can range from individually optimized determinants (e.g., Hartree–Fock states) to coherent states, enabling access to multiple eigenstates while reducing circuit depth and enhancing sampling efficiency. Notably, error mitigation techniques such as shadow tomography may also help reduce measurement overhead.
@@
Adaptive, parallel, and heuristic enhancements to VQE. Adaptive strategies like ADAPT-VQE and operator pooling dynamically construct compact ansatzes tailored to the correlation structure of the system, minimizing circuit depth without sacrificing accuracy. In parallel, recent work has introduced techniques for parallel parameter optimization to accelerate convergence in VQE [184, 185]. Heuristic methods inspired by quantum annealing have also been explored to guide variational optimization and improve initial state preparation [186]. Complementing these efforts, quantum-information-informed approaches such as PermVQE [187] and ClusterVQE [188] exploit mutual information to reduce circuit depth via entanglement-aware qubit permutation and to enable scalable parallel VQE through graph-based decomposition, respectively.
@@
Automation and user-accessible deployment of quantum solvers. Recent efforts have focused on automating the ansatz and subspace construction process to facilitate the broader adoption of non-orthogonal quantum solvers [189]. These efforts aim to make quantum algorithm deployment more accessible to non-experts by reducing the manual complexity of circuit design, subspace definition, and error mitigation integration.

### v2

@@ III.2 Quantum Dynamics and Noise-Informed Simulation
Quantum devices may be well-suited to this regime due to their natural ability to implement unitary time evolution and sample from high-dimensional entangled states. Methods such as Trotterized real-time evolution, variational dynamics [113], and Krylov-subspace propagation (related to quantum signal processing, QSP, or Qubitization) have been proposed as viable quantum algorithms (see e.g. Ref. 107 for a recent overview). However, challenges remain, including the accurate preparation of initial states (which can be non-trivial and propagate errors), significant measurement overhead due to wavefunction collapse, and the difficulty of accurate Trotterization, especially for coupled, time-dependent systems.
@@
Simulating OQS presents the additional challenge that quantum computers naturally perform unitary evolution, while open systems exhibit non-unitary dynamics due to environmental interaction. Standard quantum approaches to OQS include embedding the system within a larger environment simulated unitarily (ancilla-based methods), using stochastic quantum trajectories, or implementing Kraus operators [114, 115, 116, 117, 118, 119, 120]. Intriguingly, noise, traditionally viewed as an impediment, has been proposed as a resource [121]. Hardware-induced decoherence, if properly characterized or engineered, may serve as a proxy for environmental interactions, thereby facilitating OQS simulation. This “noise-assisted” approach has been explored in analog and digital quantum computing [122, 123, 124, 125], potentially reducing overhead compared to full error mitigation by using techniques like partial error correction [126, 127] or pulse control. From the perspective of quantum utility, the significance of treating noise as a resource is not that noise can be engineered (classical simulations can emulate noisy channels), but that hardware-induced decoherence can be harnessed in situ within quantum workflows. This enables hybrid strategies where noise directly contributes to modeling open-system dynamics at scales beyond tractable classical simulation, linking the presence of hardware imperfections to scientifically meaningful predictions.
@@
The encoding of bosonic modes, representing vibrational, solvent, or bath degrees of freedom, also requires careful consideration. Strategies include truncated Fock spaces, coherent-state encodings, and squeezed-state representations [128]. The simplified operator structure in exciton–boson models (often or terms, versus for generic electronic Hamiltonians) is advantageous for both classical and quantum solvers. In the fault-tolerant setting, it also translates directly into lower block-encoding and Trotter (or linear combination of unitaries, LCU) costs (i.e., smaller term sums and norms), fewer oracle calls under qubitization, and simpler state-preparation primitives (e.g., Gaussian/coherent/squeezed states and qumode encodings), thereby reducing the logical-qubit and -count budgets for dynamics simulations. We emphasize that classical methods benefit from this structure as well; however, long-time, finite-temperature, or non-Markovian dynamics remain challenging in practice due to tier growth in HEOM, the curse of dimensionality in MCTDH, and entanglement-growth-limiting tensor networks. By contrast, fault-tolerant quantum simulation avoids the real-time sign problem and can exploit the reduced term count to achieve gate/query complexities that scale with the Hamiltonian norm and simulation time, making exciton–boson platforms promising early targets for resource-aware quantum dynamics [110, 109, 107, 79, 80, 81]. Furthermore, embedding approaches and the development of advanced Gaussian ansatzes are being explored to reduce the resource requirements for representing these bosonic degrees of freedom [129]. Non-Hermitian Hamiltonians are also relevant for modeling dissipation, although simulating their dynamics on quantum hardware often involves mapping back to larger unitary systems or specific simulation techniques [119].
@@ III.3 Hybrid Pipelines and AI Integration
The integration of quantum computing (QC) with artificial intelligence (AI) and high-performance computing (HPC) is rapidly advancing, with major initiatives developing platform-level solutions for accelerating chemistry and materials discovery [130, 131, 132, 27, 133, 134]. AI’s role spans the entire quantum computing stack [130]. In hardware development, it accelerates qubit characterization, architecture exploration, and control pulse optimization. During operation, AI automates calibration and tuning, enabling closed-loop control strategies adaptive to evolving noise environments. In software, AI aids circuit synthesis and compression, variational optimization, and hybrid workload scheduling. It enhances QEC by improving decoder performance [32] and enabling scalable, low-latency strategies. In post-processing, AI reduces measurement overhead and mitigates errors in tasks like tomography [135], observable estimation, and readout classification. This end-to-end integration of AI is increasingly viewed as essential for making scalable, fault-tolerant quantum computing practical, particularly in the near term. AI contributes by improving decoder throughput for QEC, reducing calibration and scheduling overhead, and lowering measurement costs, which are bottlenecks that otherwise limit the effective utilization of hybrid quantum–classical resources.
Recent efforts illustrate how industry and academic platforms are beginning to integrate QC, AI, and HPC resources in practice. We cite several examples below only as representative implementations, while the key role of such infrastructures is to enable tight feedback loops, reduce measurement overheads, and improve reproducibility in early fault-tolerant quantum chemistry. One effort is NVIDIA’s DGX Quantum system, which enables low-latency, tightly coupled execution between quantum processing units (QPUs) and GPUs to support real-time AI-assisted QEC, calibration, control, and readout [130]. To program such heterogeneous quantum–classical systems, NVIDIA introduced the CUDA-Q platform [136], a single-source, hardware-agnostic programming framework that unifies quantum and classical workflows, leveraging NVIDIA’s existing CUDA and AI ecosystems. Another effort is provided by Microsoft’s Discovery platform (formerly, Azure Quantum Elements), which offers a cloud-integrated platform that provides access to multiple quantum hardware backends alongside Azure’s HPC infrastructure and AI toolkits, enabling users to build scalable hybrid quantum applications within a managed cloud environment [131]. Complementing these, IBM’s Qiskit Runtime provides a cloud-native execution model with low-latency sessions and server-side primitives (Estimator, Sampler) that streamline hybrid quantum–classical loops and support IBM’s roadmap toward tighter integration of quantum and classical resources for scientific computing [137].
In such hybrid ecosystems, AI models, which are often trained on large synthetic datasets generated via HPC simulations or augmented with quantum data, are used to accelerate discovery by enabling rapid screening, property prediction, and system-level optimization. For example, AI models deployed within Azure Quantum Elements have been used to evaluate millions of potential battery materials [138], while NVIDIA has developed generative AI approaches such as GQE [139] and QAOA-GPT [140] to assist in synthesizing quantum circuits with features such as reduced depth and enhanced expressivity. In addition to algorithm development, NVIDIA has also demonstrated the use of AI for QEC in collaboration with QuEra [141].
This tiered, adaptive computational approach allocates quantum resources to the most challenging subproblems where classical methods struggle, such as the transition-state barrier heights poorly captured by DFT or systems with strong multireference character [142, 143]. AI is integral across this hybrid workflow: from hardware design and calibration to device control, algorithm optimization, QEC decoding, and post-processing [144, 145]. Diagnostic tools, including uncertainty quantification in AI models, convergence analysis of classical solvers (e.g., DMRG), and AI-driven screening and circuit optimization, support dynamic resource allocation across the classical-quantum-AI stack [146, 147, 148].
@@ III.4 Benchmarking and Experimental Validation
Robust benchmarking remains essential for validating quantum methods and ensuring reproducibility. A three-way validation framework, incorporating quantum simulations, high-level classical reference data, and experimental observations, has been endorsed [49, 50, 21]. Complementary to these strategies, stabilizer circuits, comprising Clifford-only operations, offer a powerful route for algorithmic validation. Because they can be simulated efficiently on classical hardware [149, 42, 32, 43], they enable cross-checking of quantum compilation, execution, and error correction pipelines at a scale comparable to chemically relevant circuits. Embedding stabilizer-based benchmarks within hybrid workflows thus provides a practical way to verify algorithmic performance in the 25–100 logical-qubit regime before deploying more general, classically intractable simulations. This framework allows for iterative refinement of algorithms and facilitates the identification of performance gaps.
Validation when classical ground truth is unavailable. For systems that are classically intractable, we advocate a three-tier validation ladder [20, 21, 27]: (i) algorithmic self-checks, enforcing symmetries and conservation laws, variational bounds where applicable, Hellmann–Feynman consistency between energy derivatives and forces, and sum-rule/Kramers–Kronig-type constraints for response functions; (ii) reduced/embedded cross-checks, validating predictions on classically tractable fragments or limits of the model (e.g., weak-coupling, non-interacting, or localized-fragment regimes), and bracketing predictions with multiple physically motivated embeddings/downfoldings; (iii) experiment-linked, environment-aware validation, including comparison to observables measured under well-documented conditions, emphasizing differential quantities (e.g., isotope shifts, redox or excitation differences within a series, spectral shifts upon functionalization) that are less sensitive to absolute environment errors. For embedding workflows, we recommend reporting embedding-specific convergence diagnostics, such as chemical-potential and density residuals, fragment–bath mismatch, and sensitivity to solver noise, together with variance estimates if the fragment solver is stochastic. When quantum solvers are used, stabilizer-circuit checkpoints provide classically verifiable validation of compilation and execution paths before running classically intractable circuits.
Classical convergence scaffolding. Many components of hybrid quantum–classical workflows can be benchmarked purely classically prior to any quantum execution and should be used to define, justify, and bound the target instances [53, 55, 72, 74]. We recommend: (i) active-space growth along entanglement/occupation diagnostics (e.g., natural-orbital occupations, single-orbital entropies) with observables plotted vs. active-space size; (ii) embedding region/bath enlargement with convergence and stability checks (e.g., DMET chemical-potential/correlation-potential residuals, fragment–bath mismatch) under solver noise; (iii) downfolding/renormalization controls (commutator truncation order in SES/DUCC, and tensor-factorization ranks swept with observables vs. control parameter; and (iv) for dynamical/open-system models, bath discretization or tier-depth and mode truncation sweeps to establish time/frequency window fidelity. For each sweep, report stopping tolerances, monotone trends where expected, and confidence intervals; release the raw data/plots as part of the provenance package to support reproducibility.
To support both capability validation and scientific challenge, we organize benchmarks into infrastructure (Class 0) and challenge (Class 1/2) tracks. Class 0 benchmarks comprise chemically grounded but classically tractable sets (e.g., QM9/QH9, MultiXC-QM9, VQM24) [150, 151, 152, 30] that enable cross-stack reproducibility, cost-vector calibration, and measurement-reduction studies, often alongside stabilizer-circuit checkpoints. Here, the cost-vector refers to the set of quantitative resource requirements reported for a given simulation, including logical-qubit counts, depth/-count, post-downfolding term counts/ norms, and shot budgets needed to achieve target precision. This compact representation enables standardized comparisons across algorithms, hardware platforms, and benchmark instances. Class 1/2 benchmarks are also chemically relevant and experimentally tractable, but target computationally demanding instances with multireference character, coupled dynamics, or environment effects (following Ref. 153; see also Sec. IV.5). Specifically, we categorize Class 1/2 benchmarks as follows,
These systems allow simulations at different levels and experiments to interact closely, helping guide method development, hardware requirements, and software stacks across multiple disciplines.
Table 1 reflects this classification, where Class 0 entries are labeled with , and identifies several model systems as focal points for algorithm validation and workflow integration. These span catalysis, photochemistry, and energy materials. For reproducibility, we recommend releasing provenance-rich artifacts (input decks, circuits/Hamiltonians, calibration snapshots, decoder outputs, ensemble definitions, and post-processing scripts) to enable independent audit of environment assumptions and uncertainty budgets. As discussed in Section III.A, while many small active spaces remain tractable with classical solvers, larger chemically meaningful fragments often require quantum resources; the benchmarks here target such cases. To illustrate this connection, Table 2 provides resource estimates (logical qubits, circuit depth, and shot counts) for selected exemplar systems, linking orbital counts of practical active spaces to the 25–100 logical-qubit regime. As seen in Table 2, systems such as Fe2S2 clusters or small chromophores map naturally onto the 30–40 logical-qubit range when treated via embedding or downfolding, but still demand circuit depths on the order of and shot budgets exceeding . These quantitative estimates highlight the centrality of measurement-efficient ansatzes (Sec. IV.1) and error-aware phase estimation variants (Sec. IV.3) to realize practical utility in this regime.
At the same time, it is important to recognize that even when a solver achieves “chemical accuracy” on a model Hamiltonian, this does not necessarily translate into agreement with experiment. Discrepancies can arise from missing environmental effects, conformational sampling, or uncertainties in experimental reference data. Therefore, the community should view chemical accuracy as a baseline threshold, while setting the bar for success in terms of whether quantum simulations can deliver systematic, improvable accuracy for observables of chemical interest (e.g., excitation energies, redox potentials, rate constants). In this context, benchmarking strategies should emphasize not only total energies, but also chemically relevant quantities that can be cross-validated with both experiment and high-level classical methods.
Environment-aware protocol. When environment effects dominate uncertainty, the validation target should be defined as an ensemble average over conformers/solvation states [54, 58]. Concretely: (a) specify how the ensemble is generated (e.g., molecular dynamics/Monte Carlo sampling, docking, or curated conformer sets), how weights are assigned (e.g., Boltzmann, reweighting), and how observables are averaged; (b) report a sensitivity analysis to environment modeling choices (continuum vs. explicit solvent, ionic strength, counterions, pH), and propagate these to an uncertainty budget alongside quantum-solver statistical/systematic errors; (c) prioritize differential observables and provide stated error bars that reflect both ensemble and hardware/runtime contributions. When quantum subroutines (i.e., the quantum-executed portions of a hybrid workflow) are used, include stabilizer-circuit check-points as classically verifiable guards within the workflow.
For example, while absolute redox potentials of Fe–S clusters in solution or protein environments are far beyond current quantum resources, simplified embedded fragments (e.g., Fe2S2, Fe4S4) can serve as tractable benchmarks, with solvation and environmental contributions incorporated through classical embedding or hierarchical workflows.
In summary, identifying chemically meaningful opportunities and establishing robust validation protocols highlight the need for carefully chosen algorithmic strategies. We turn to these strategies in Sec. IV, focusing on approaches designed to operate effectively within the 25–100 logical-qubit regime.
| Model/System | Active Space (orbitals) | # of Logical Qubits | Circuit Depth/-count | # of Shots | |
| black Fe2S2 cluster (bioinorganic model) | orbitals (embedded) | 30–50 | – | – | |
| Chromophore (photochemistry) | 20–30 orbitals | 25–30 | – | – | |
| Benzene (aromatic benchmark) | 30 orbitals | 35–45 | |||
| Alkali hydrides (NaH, KH) | 10–12 orbitals | 20–25 |
@@ IV Algorithmic Innovations
In this section, we survey algorithmic strategies most relevant for the 25–100 logical-qubit regime, focusing on representative methods, their theoretical kernels, resource requirements, and trade-offs with classical approaches.
@@
The development of quantum algorithms capable of addressing chemically relevant problems using 25–100 logical qubits must be guided by co-design principles that account for fault-tolerant constraints. These include limited circuit depth, restricted qubit connectivity, noise resilience, and modular architecture [172, 173]. The goal is to create a unified framework for evaluating algorithms across diverse problems, establishing robust metrics and methodologies to uniformly measure performance [174]. Rather than converging on a single dominant paradigm, the current landscape favors the parallel exploration of diverse algorithmic strategies tailored to specific problem structures and hardware capabilities [144, 20, 21]. Many algorithms exist, balancing accuracy and efficiency, as exact solutions scale exponentially and are often impractical for large systems.
Scope and selection rationale. The directions highlighted below were selected using criteria tailored to the 25–100 logical-qubit regime: (i) ability to meet tight logical-qubit, depth, and -count budgets (including modest ancilla and code distance); (ii) alignment with chemistry workflows (active-space/embedding, downfolding, and dynamics) with observable-level outputs; (iii) mitigations for noise and trainability (e.g., shallow/subspace circuits, iterative/filtered QPE, error-aware outer loops); (iv) transparent resource models and benchmarkability; and (v) a clear path toward fault-tolerant realization. Methods that currently depend on resources outside this envelope (e.g., deep phase estimation without compression, qRAM-intensive routines) were deprioritized for this near-term discussion. Given known impediments in variational quantum algorithms (VQAs), such as the measurement overhead, barren plateaus, gate-fidelity thresholds, and classical simulability in plateau-free regimes [175, 176, 177], we include only shallow/subspace variational approaches paired with measurement-reduction and noise-aware outer loops; hardware-agnostic, deep VQE heuristics are deprioritized in this regime. As summarized in Fig. 2, these directions must be understood in light of the contrasting scaling behaviors of classical and quantum approaches. In contrast, resource-aware quantum algorithms such as qubitization-based QPE or measurement-efficient VQE variants target lower-order polynomial gate/query complexity, albeit with explicit shot budgets. Within the 25–100 logical-qubit band, this yields the conservative but realistic ordering that FCI remains several orders of magnitude more costly than Selected CI/DMRG, which in turn is about two orders higher than resource-aware quantum approaches. This hierarchy motivates the structured ansatzes, efficient measurement strategies, downfolding and dimensional-reduction methods, and compressed QPE techniques discussed in the following subsections.
IV.1 Structured Ansatzes and Measurement-efficient Subspaces
Why included: shallow/subspace constructions offer multi-state access within strict depth/ budgets, integrate with downfolded or embedded active spaces, and can be paired with measurement-reduction (e.g., shadows/overlap estimators) and noise-aware outer loops for trainability.
Variational quantum algorithms (VQAs), including VQE [65], face several well-documented challenges. Among these, shot noise and measurement overhead are often cited as particularly demanding in practice, since sample requirements scale with Hamiltonian variance and target precision. However, even in noise-free emulators, VQE remains a heuristic constrained optimization without guaranteed convergence, and practical performance depends strongly on the ansatz, optimizer, and problem structure [29]. These limitations, together with barren plateaus, spin contamination, and hardware noise, underscore why structured ansatzes and measurement-efficient strategies are essential. At the same time, rapid progress in hardware [178, 179, 48] provides a dynamic backdrop in which these challenges must be continually re-evaluated. While neural network ansatzes have been explored, particularly for solids and lattice models [180], they are not chemically intuitive and remain less common in quantum chemistry. By contrast, most innovation for molecular applications has focused on chemically inspired and hardware-efficient ansatzes [35, 181]. We have identified four major directions below.
@@ IV.1 Structured Ansatzes and Measurement-efficient Subspaces
First, different parameterizations of the unitary Cluster Jastrow (uCJ) ansatz [182, 183] have been explored. In the original formulation [182], the orbital-rotation generator is a complex anti-Hermitian matrix, which in practice is often restricted to a real antisymmetric form, while the Jastrow matrices are pure imaginary and symmetric. This “real / imaginary ” restriction is the most widely used, as it yields a compact, chemically meaningful parameterization that captures strong correlation more effectively than generalized unitary coupled cluster with singles and doubles (GUCCSD) [184], while significantly reducing quantum resource requirements [183]. Alternative parameterizations such as the imaginary symmetric (Im-uCJ) and general complex (g-uCJ) forms are also promising, particularly for bond dissociation. These findings suggest that tuning the parameterization strategy can improve accuracy and robustness, though care must be taken to evaluate potential drawbacks such as spin contamination. We treat adaptive/heuristic enhancements (e.g., operator pooling, layer-wise updates) as optional refinements, contingent on reporting trainability diagnostics and measurement budgets for target observables. The measurement cost of VQE approaches can be quantified as
where is the number of measurements required, is the variance of the Hamiltonian, and is the target precision. This scaling highlights why measurement-reduction strategies (e.g., classical shadows, low-rank factorizations, grouping of commuting terms) are critical for pushing VQE into the 25–100 logical-qubit regime.
@@
Second, optimization-free and adjustable subspace construction methods provide versatile alternatives. These include density-matrix-based quantum subspace diagonalization approaches such as quantum subspace expansion (QSE) [185], as well as Krylov/Lanczos methods [186, 187, 188, 189], non-orthogonal configuration interaction schemes [172], quantum computed moment techniques [190], and generator coordinate-inspired strategies [191]. The subspace basis can range from optimized determinants (e.g., Hartree–Fock states) to coherent states, enabling access to multiple eigenstates while reducing circuit depth and enhancing sampling efficiency. Error mitigation techniques such as shadow tomography [192, 193] may further reduce measurement overhead. Representative resource estimates and trade-offs for these methods are illustrated in Deep-Dive IV.1.
@@
Third, adaptive, parallel, and heuristic enhancements to VQE aim to address trainability and efficiency. Adaptive strategies such as ADAPT-VQE [194, 195] and operator pooling dynamically construct compact ansatzes tailored to the correlation structure, minimizing depth without sacrificing accuracy. Parallel parameter optimization has been introduced to accelerate convergence [196, 197], and heuristic methods inspired by quantum annealing have been applied to improve state preparation and guide optimization [198]. Complementary approaches like PermVQE [199] and ClusterVQE [200] use entanglement-informed qubit permutations or graph-based decompositions to reduce depth and enable scalable parallelization.
@@
Finally, automation and user-accessible deployment of quantum solvers are gaining traction. Recent efforts focus on automating ansatz and subspace construction to facilitate broader adoption of non-orthogonal quantum solvers [201], thereby reducing the manual complexity of circuit design, subspace definition, and error mitigation integration. Such developments aim to make quantum solvers accessible to a wider community beyond quantum algorithm specialists. For clarity, the main families discussed above are organized in Table 3, which summarizes kernels, resource profiles, and trade-offs in the 25–100 logical-qubit regime.
Taken together, these structured ansatzes and subspace strategies illustrate the trade-off between expressivity and resource efficiency. The common theme across uCJ, Krylov/subspace, and adaptive refinements is to leverage chemically motivated structure, shallow depth, and measurement reduction to achieve tractable scaling under 25–100 logical-qubit budgets. These methods serve as modular building blocks for embedding and downfolding frameworks (Sec. IV.2), where compact effective Hamiltonians can be paired with such solvers.
| Method | Key Idea / Kernel | Notes for 25–100 logical-qubit regime | |||||||||||
| Heuristic refinements |